40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 12 20:15:22 2016, Lydia, Update, SUS, Diagonalization in air ITMY_fit.jpgBS_diag.jpg
    Reply  Tue Sep 13 19:18:43 2016, Lydia, Update, SUS, Diagonalization in air ITMY_diag.jpg38.png
       Reply  Wed Sep 14 20:27:03 2016, Lydia, Update, SUS, Diagonalization ETMY_diag.png
          Reply  Thu Sep 15 18:37:20 2016, Lydia, Update, SUS, Diagonalization ETMY_osemspec.pngITMY_osemspec.png
             Reply  Fri Sep 16 19:14:27 2016, Lydia, Update, SUS, Diagonalization ETMY_yawdrivedefault.pngETMY_yawdrivenew.pngETMY_yawdriveold.png57.png
Message ID: 12495     Entry time: Wed Sep 14 20:27:03 2016     In reply to: 12490     Reply to this: 12497
Author: Lydia 
Type: Update 
Category: SUS 
Subject: Diagonalization 

Today the main optics were free swinging for several hours, so I attempted diagonalization in vacuum.

  • ITMY still has bad phases. I looked at the spectra for this and other optics, and it looks like the other optics have the 60Hz line notched out for all coils while ITMY only has it notched on the side coil. (Using C1:SUS-ITMY_SENSOR channels). Where is this controlled from, and could it be the source of the issue? 
    • I tried using a different coil as the "standard," with the other coils compared against it in tfestimate. Default is UL, I tried UR and LL. The phase problems were still present for ITMY, but the script was still working fine for other optics.
    • The phase difference between coils is different for different start times.
    • A short segment of the time series for ITMY shows significantly more high frequency noise than for other optics at the same time.
  • The ETMY matrix for vacuum has the wrong sign for UL coupling to pitch! The diagonalization results look OK on the graph, but the butterfly mode still has small peaks (See attachment 1). When the individual coil spectra are plotted, the angular degrees of freedom show very weak coupling for UL to pitch, and LL to yaw. We initially replaced the matrix on the MEDM screen with the one generated by the script. After realizing this, the PIT row was changed to 1 1 -1 -1 0, but the effectiveness of the damping on the locked transmission fluctuations was about the same both ways.
Attachment 1: ETMY_diag.png  47 kB  | Hide | Hide all
ETMY_diag.png
ELOG V3.1.3-